### v2

# Walking the Tightrope of LLMs for Software Development: A Practitioners’ Perspective

Abstract
Background: Large Language Models emerged with the potential of provoking a revolution in software development (e.g., automating processes, workforce transformation). Although studies have started to investigate the perceived impact of LLMs for software development, there is a need for empirical studies to comprehend how to balance forward and backward effects of using LLMs. Objective: We investigated how LLMs impact software development and how to manage the impact from a software developer’s perspective. Method: We conducted 22 interviews with software practitioners across 3 rounds of data collection and analysis, between October (2024) and September (2025). We employed Socio-Technical Grounded Theory for Data Analysis (STGT4DA) to rigorously analyse interview participants’ responses. Results: We identified the benefits (e.g., maintain developer flow, improve developer mental models, and foster entrepreneurship) and challenges (e.g., damage to developers’ reputation) of using LLMs at individual, team, organisation, and society levels; as well as actionable guidances into how mitigate these challenges. Conclusion: Critically, we present the trade-offs that software practitioners, teams, and organisations face in working with LLMs. Our findings are particularly useful for software team leaders and IT managers to assess the viability of LLMs within their specific context.
Index Terms:
Software Engineering, Artificial Intelligence, Large Language Models, Socio-Technical Grounded Theory, Interviews
1 Introduction
“There is nothing permanent except change.”Heraclitus
This quote by the Greek philosopher Heraclitus highlights how the world is continuously changing. Large Language Models (LLMs) are the contemporary catalyst for a revolution in the Information Technology sector [45], starting from the release of LLM tools like ChatGPT and GitHub Copilot for the general public between 2022 and 2023 [55, 15]. LLM-powered code generators and assistants like GitHub Copilot, for instance, fostered the emergence of a new potential pillar for software development: AI pair programming [10, 58]. LLMs can support a variety of software development tasks, such as code generation and information retrieval [24, 57, 13]. Enterprise LLM adoption reports from McKinsey [40] and DORA [11, 12] call attention to growing interest from companies in examining the potential of LLMs for software development. There is also a growing shift from traditional Q&A online communities like Stack Overflow towards LLM tools as the first source of support [45, 6]. For instance, the decline in networking traffic of Stack Overflow may be attributed to developers adopting LLMs [27].
Many investigations (e.g., [59, 9, 16, 43]) highlight the potential benefits associated with software practitioners adopting LLMs. For instance, Cui et al. [9] found that software developers using GitHub Copilot achieved an increase of 26.08% in the number of weekly completed tasks when conducting an experiment with almost five thousand software developers from companies including Microsoft and Accenture. On the other hand, many studies (e.g., [31, 30, 32, 8]) present the downsides of using LLMs. For instance, Lee et al. [32] observed from surveying 319 knowledge workers that LLMs can potentially affect software developers’ critical thinking skills.
Although there is an emerging amount of studies related to LLMs in Software Engineering (SE) [24, 50], there is still a need for investigations focusing on managing the impact of LLMs. Mohamed et al. [42] conducted a systematic literature review on how LLMs affect software developer productivity. From their thirty-seven selected studies, they summarised most benefits and risks concerning how LLMs affect software developer productivity, such as supporting knowledge acquisition and promoting over-reliance, disruptions to developer flow. They highlight that few studies explore aspects involving communication and human-human collaboration. Ferino et al. [19] conducted a systematic review exploring novice software developers’ adoption and use of LLMs in SE activities. From their 80 selected studies, they identified many research gaps, such as exploring the impact of LLMs on mentorship interactions.
To gain more understanding about the impact of LLMs in SE tasks and especially how to manage it, our investigation focused on answering this main question: How does the use of LLMs for software development impact software practitioners? This question was decomposed into the following research questions (RQs):
  • RQ1. How do LLMs help software developers move forward? – Moving developers forward involves the experienced benefits of using LLMs.
  • RQ2. How do LLMs create an imbalance in software development? – Creating an imbalance involves experienced challenges when using LLMs. We draw special attention to this RQ in this report, as its findings may guide future studies.
  • RQ3. How can software developers achieve a balanced use of LLMs? – It encompasses actionable recommendations to use LLMs.
To answer these RQs, especially RQ2, we conducted twenty-two semi-structured interviews with software practitioners across three rounds - between October (2024) and September (2025). Our study aims to comprehend the current industrial perspective on software developers adopting Large Language Model-based tools on SE-related activities, which involves exploring the benefits, challenges, limitations, and recommendations shared by software practitioners involved in SE-related activities. Our analysis reveals the following main benefits: (B1) reduced effort due to LLMs as a foundation to boost code development and perceived saving time; (B2) flow experience when LLMs mitigate interruptions and automate simple and tedious tasks; (B3) interaction with LLMs influencing developers’ personality and creating a safe space; and (B4) LLMs promoting LLM entrepreneurship as a consultant for every (not complex) question; and these challenges: (D1) developer flow interruptions due to frictions in human-LLM interaction; (D2) reduced effort due to LLMs eroding developer skills and agency; (D3) interaction with LLMs reducing mentorship opportunities. The main contributions of this research include:
  • Identification and categorisation of benefits and challenges of adopting LLMs for software development tasks in terms of individual, team, organisation, and society level.
  • A set of recommendations on how to best use LLMs for software development tasks.
2 Methodology
Figure 1: Iterative steps of data collection and analysis.
This study aims to comprehend the current industrial perspective on the impact of software practitioners adopting LLM-based tools for SE-related activities. We conducted an interview-based study with 22 software practitioners across three rounds of data collection [^1] Approved by Monash Human Research Ethics Committee (No. 44875). We applied socio-technical grounded theory for data analysis (STGT4DA) [22] to analyse the interview data iteratively. In our first round, we conducted six interviews with novice software practitioners. In our second round, we conducted interviews with thirteen novice (n=10) and experienced (n=3) software practitioners, resulting in the emergence of concepts and categories. Finally, we interviewed three more software practitioners during the third round of data collection and analysis, with the aim of consolidating our findings. Figure 1 provides an overview of our study methodology. The supplementary material, which includes the pre-interview questionnaire, interview guide, STGT4DA coding example, assumption list, and LLM capabilities, is available online [18].
TABLE I: Participant Demographics.
ID Role Domain Country of Residence Total Experience Self-Reported Level of Experience Round
P1 Software Developer IT Brazil 3-5 Years Experienced 1
P2 Software Engineer IT Australia 1-2 Years Experienced 1
P3 Software Engineer IT Australia 3-5 Years Intermediate 1
P4 Software Developer Finance Brazil 3-5 Years Intermediate 1
P5 Software Developer IT Australia 0-0.9 Years Intermediate 1
P6 Data Engineer DS/BD Brazil 3-5 Years Highly Experienced 1
P7 BI Analyst TELCOM Australia 0-0.9 Years Intermediate 2
P8 Software Engineer IT Australia 1-2 Years Experienced 2
P9 AI Engineer Government Brazil +10 Years Intermediate 2
P10 Software Engineer IT Canada 6-10 Years Highly Experienced 2
P11 Web Developer IT Australia 3-5 Years Highly Experienced 2
P12 Software Engineer Healthcare United States 6-10 Years Experienced 2
P13 Software Developer IT Australia 1-2 Years Advanced Beginner 2
P14 Data Engineer TELCOM Australia 1-2 Years Intermediate 2
P15 Researcher Engineering Australia 0-0.9 Years Novice 2
P16 R&D Researcher Gaming Canada 1-2 Years Intermediate 2
P17 Data Analyst IT Australia 1-2 Years Intermediate 2
P18 Software Developer Finance Finland 3-5 Years Intermediate 2
P19 ML Scientist IT Canada 3-5 Years Experienced 2
P20 Software Engineer IT Malaysia 1-2 Years Intermediate 3
P21 Research Engineer IT Singapore 1-2 Years Experienced 3
P22 Software Developer IT United States +10 Years Highly Experienced 3
IT: Information Technology; DS/BD: Data Science/Big Data; TELCOM: Telecommunications.
2.1 Study Design and Piloting
Based on the findings (i.e., benefits, challenges, recommendations) from our systematic literature review [19], we developed a preliminary interview guide, which also included potential follow-up questions. This interview guide [18] was improved based on discussions with the PhD supervisors, as well as feedback from an industrial collaborator, who has experience managing a software team in an Australian company. The Attitude Towards Artificial Intelligence (ATAI) scale, developed by Sindermann et al. [51], was included in the pre-interview questionnaire to provide an initial understanding of participants’ attitude towards AI, allowing the interviewer (first author) to adjust interview questions based on each participant’s strongest responses across five dimensions — perceived benefits, trust, fear, job displacement, and existential concerns — and also afforded additional contextual information to support the subsequent data analysis (i.e., open coding and memoing). ATAI scale includes the following five nine-point scale questions, which we adapted to a five-point scale ranging from strongly disagree to strongly agree to facilitate the participants to answer them:
  • I fear artificial intelligence
  • I trust artificial intelligence
  • Artificial intelligence will destroy humankind
  • Artificial intelligence will benefit humankind
  • Artificial intelligence will cause many job losses
From our industry collaborator’s feedback, we included questions, for example, seeking to understand participants’ perceptions of the impact of LLMs on their career trajectory or job market competitiveness in the near future. We collected participants’ information regarding basic demographics, work experience, and experience with LLM tools. We also conducted two pilot interview sessions with experienced software practitioners to evaluate the interview structure - whether to adopt a pre-interview questionnaire to collect participants’ information - and the clarity of the interview questions. During our pilot study, we observed that it takes only about 5-10 minutes to collect participants’ information during the interview; however, we also observed that the pilot study participant who filled out a pre-interview questionnaire was more comfortable (relaxed) during the interview. We believe that while using the pre-interview questionnaire, with the interview moment focused only on the interview guide, we provided a more simplified structure for the participant. The pilot interview using a pre-interview questionnaire was not included in the analysis because this participant had no experience using LLMs for software development.
2.2 Sampling
Leveraging the findings from our SLR [19] on novice developers’ adoption of LLMs for SE tasks as a foundation, we began with purposive sampling in the first round, focusing on less experienced developers with less than five years of professional experience (See Fig. 1). In the following rounds, we moved to convenience sampling, interviewing software practitioners from all levels of experience since our first round of data collection and analysis did not result in enough codes, concepts, or memos to justify continuing only with novice participants. This would also support further comparison between novice and experienced developers’ perceptions and experiences. We advertised this study on our professional social media, LinkedIn and X (formerly Twitter). We also recruited participants from our personal connections. Initially, we advertised our study, including a \$50 (AUD) voucher gift card. However, due to potential fake participants identified based on discrepancies between their IP addresses and the countries they submitted in the pre-interview questionnaire, we got approval from our Faculty Ethical Review Committee to omit the gift card in the advertisement, and only offered it to genuine participants at the end of the interview. We also adopted a snowballing approach by encouraging the participants to share our study with others.
TABLE II: Participants’ use of LLMs.
Exp. (LLMs in General) # of Practitioners Exp. (LLM4SE) # of Practitioners # of Prompts per Day # of Practitioners
Less than 1 year   5 Less than 6 months   4 Less than or equal to 2   2
Less than 1.5 year   5 Less than 1 year   8 3 - 5   5
Less than 2 years   7 Less than 1.5 year   3 6 - 10   3
More than 2 years   5 Less than 2 years   5 11 - 20; 21 - 30   4
More than 2 years   2 50 - 100; +100   2
ATAI Scale Frequency* SE Tasks supported by LLMs # of Practitioners
I fear artificial intelligence       [7, 7, 4, 2, 2] Code-related tasks   20
I trust artificial intelligence      [4, 10, 4, 4, 0] Test-related tasks   11
Artificial intelligence will destroy humankind       [11, 3, 3, 4, 1] Documentation-related tasks   9
Artificial intelligence will benefit humankind      [1, 0, 3, 12, 6] Requirement-related tasks   7
Artificial intelligence will cause many job losses      [0, 2, 7, 10, 3]
LLM tools # of Practitioners LLM tools # of Practitioners Company’s AI Policy # of Practitioners
ChatGPT   18 Cursor   2 Allowed to use   16
GitHub Copilot   7 Phind   2 No policy   3
Claude   7 Mistral   2 Prohibited to use   1
Llama   4 Jan.AI   1 I do not know   1
Gemini   2 H2O Danube   1 Prefer not to say   1
Perplexity   2 Continue   1
Microsoft Copilot   2
*Order of the bars in Frequency column of ATAI Scale graph: Strongly disagree → Somewhat disagree → Neither agree nor disagree → Somewhat agree → Strongly agree, followed by respective values.
The first round of interviews with six novice software practitioners took place between October and November (2024). Then, after changes in the interview guide, the second round of interviews with thirteen software practitioners took place between April and June (2025). The third round of interviews occurred in September (2025). The interviews were conducted and recorded via Zoom meetings, which also provided the transcriptions. Interviews were scheduled to go 40-45 minutes, but ranged from 34 to 58 minutes. We manually reviewed the transcriptions, filtered any sensitive information, and de-identified the participants.
Participants’ demographics. Our participant pool covers a diverse range of roles, countries of residence, years of experience, and experience using LLMs. We conducted twenty-two interviews with software practitioners from Latin America, Oceania, North America, Asia, and Europe - mostly from Australia (n=10). Table I provides an overview of participants’ demographics. Our study participants worked in different domains, such as healthcare, government, and finance. Most of the participants are males (n=14), have equal to or less than five years of professional experience (n=13), and describe their skill level as intermediate or advanced beginner (n=17).
Regarding their experience using LLM tools, which can influence the faced challenges and suggested solution strategies, participants reported using a variety of LLM tools. Not surprisingly, ChatGPT (n=18) was the most recurrent tool used by the interview participants. Table II summarises participants’ demographics information related to the adoption of LLM tools. Most of the participants (n=14) have used LLMs in general, not specifically for SE tasks, for more than one year. According to their self-report amount of daily prompts, 59% (n=13) of participants engage daily with LLMs through at least 6 prompts. Their attitudes towards LLMs captured via the Attitude Towards Artificial Intelligence (ATAI) scale highlight different inclinations towards AI technologies. While mostly (n=14) strongly disagreeing or somewhat disagreeing about fearing AI, they also mostly (n=14) disagree or somewhat disagree about trusting AI. Most (n=18) somewhat agree or strongly agree that AI will benefit humanity, and somewhat disagree or strongly disagree that AI will destroy humanity (n=14). However, most participants (n=13) somewhat agree or strongly agree that AI will affect the job market, leading to job losses. With respect to their experience using LLMs for SE tasks, mostly (n=16) reported having more than six months of experience, and they use LLMs mostly for code-related tasks (n=17), such as coding and debugging. Most of the study participants (n=14) reported that their companies allow them to employ LLM tools for job responsibilities.
The fast-paced evolution of LLM tools compels us to examine the features available in the tools mentioned by interview participants. During our three rounds of data collection and analysis, we observed an emergence and evolution of different LLM tools (e.g., ChatGPT, Cursor) and integration with traditional tools (e.g., ChatGPT for Databricks [14], Copilot for Power BI [54]). This was also highlighted by the participants, e.g.: “I’ve been able to see a bit of how those tools evolve during this time. [For example, GitHub] Copilot evolved from being a better autocomplete to having more tools [and features].” – P10 [Software Engineer]. Ferino et al. [19] suggest that researchers investigating LLMs for SE should provide a snapshot of the LLM features with the intent of improving clarity of the context related to the data collection. We summarise the main features in the supplementary online package [18].
Figure 2: Emergence of the category “Impact on using LLMs” from raw datacodesconceptssubcategoriescategory through constant comparison [23].
2.3 Data Analysis
Applying STGT’s basic stage of data collection and analysis, we conducted a lean literature review to identify research gaps, followed by data collection (e.g., survey), and hashtag open coding, constant comparison, and basic memoing for data analysis [22, 23]. STGT4DA has been applied in various SE studies to identify “important patterns in the qualitative data and present them as layered and/or multi-dimensional findings, along with insights and reflections” [46, 36, 37, 21]. We leveraged its foundational socio-technical research framework to investigate the socio-technical phenomenon of software practitioners’ perspectives on the role of LLM tools in software development, in the socio-technical domain of LLMs for SE (LLM4SE), with socio-technical actors such as software developers, software engineers, data engineers, data scientists, and DevOps engineers. As socio-technical researchers, we combined our skills and experiences. The interviews were conducted by an early-career researcher (first author) under the guidance of three experienced researchers, and the data were analysed by the first author with detailed feedback from the second author. We also applied socio-technical data, tools, and techniques such as a Qualtrics survey to collect participant demographics, Zoom for recording and transcription, DeepL222www.deepl.com for translation of the interview transcriptions in Portuguese (P1, P4, P6) to English, and Nvivo and spreadsheets to support the analysis of the interview transcriptions.
Hashtag Open Coding. The process of hashtag open coding [23, Chapter 10] was iteratively improved by the first researcher following the guidance of the second author. Initially, the first author printed out one of the interviews and coded it, followed by a review and discussions with an experienced researcher. After this, the researcher coded two more interviews using Google Docs. The researcher decided to move to NVivo, instigated by its features, and spreadsheets. Before performing the hashtag open coding, and between rounds of data collection and analysis, the first author was advised to create and update an assumption list [18]. This helped to make the researcher’s assumptions and biases transparent so they could be managed.
Constant Comparison. Similar codes were grouped into concepts, and similar concepts under sub-categories, and similar sub-categories under categories. During this process, we drew diagrams to facilitate visualisation and insights regarding the relationship between concepts and sub-categories, and also conducted discussions between the authors. Figure 2 illustrates the emergence of the category Impact of using LLMs via coding and constant comparison.
Memoing. While performing constant comparison, we also wrote down memos reflecting our ideas, thoughts, and key elements emerging throughout the analysis. Those memos, which encompass both interview and inter-interview levels, supported the update of the interview guide between rounds. At the interview level, we compared the participants’ pre-interview questionnaire responses with their interview responses, especially concerning the ATAI scale. That information provided more context for our analysis. At the inter-interview level, we compared codes from different interview participants, which also helped to refine our concepts and sub-categories. An example of a memo “How do LLMs affect developer intuition?”, is shown below. The discussion on the main insights from memoing is presented in Section 4.2.
Memo on “How do LLMs affect developer intuition?” LLMs seem to emerge as a learning tool for software developers based on participants P12, P14, and P15, e.g., relying on LLMs to cover gaps in knowledge. At the same time, P14 argues that while using LLMs for learning, novice developers should do the majority of the work. This practice seems to be necessary to improve novice developers’ skills, but also to cultivate the intuition inherent in experienced developers. Intuition is something that comes naturally after fully understanding something. That intuition (internalised knowledge) helps to avoid scenarios where LLMs are misleading, which P11 described as ChatGPT being “a very good liar”.
3 Findings
We found variations in the impact of software developers adopting LLMs across four levels:
  • Impact on the Individual Level: it refers to LLMs impacting software practitioners directly;
  • Impact on the Team Level: it involves how LLMs influence software development teams and their collaboration;
  • Impact on the Organisation Level: it encompasses how LLMs affect entire software organisations;
  • Impact on the Society Level: it refers to how LLMs affect daily life in society, not just developer communities.
Tables 3.1.1 and LABEL:tab:disadvantageIndividualLevel illustrate the benefits and challenges at the individual level of using LLM tools to support software development tasks, along with the reasons influencing these impacts. For instance, Table LABEL:tab:disadvantageIndividualLevel shows that using LLMs for brainstorming may hinder developers’ skills due to a lack of practice in software development skills and overreliance on LLMs. Additionally, we have summarised the main benefits and challenges in Figures 3 and LABEL:fig:MainDisadvantagesIndividualLevel, respectively. For instance, using LLMs for debugging can boost code development, saving development time, reducing software developers’ effort, and providing learning opportunities to software developers. We organised the remainder of this section by describing thebenefits of using LLMs (section 3.1), challenges when using LLMs (section 3.2), and recommendations on how to use LLMs (section LABEL:subsec:recommendations). Throughout this section, we provide representative quotes to illustrate participants’ experiences and viewpoints. In order to protect the privacy of our participants, we refer to the interview participants as P1-P22. Similar to Masood et al.[38], we will use some expressions to indicate the extent of the predominance. In this sense, ‘few’ refers to less than 25%, ‘many or majority’ to over 50%. Although this is not a quantitative study, it indicates how much evidence supports each theme.
3.1 RQ1. How do LLMs help software developers move forward?
3.1.1 Individual Level
Reducing developer effort with LLMs. A majority of participants (i.e., P1-P18, P20-P22) described scenarios where LLMs work to lighten their efforts for simple, repetitive, or tedious tasks. This happens because software practitioners hand over the control to LLMs’ automation capabilities, e.g.: “As I have this extension in Databricks, it makes it much easier. For example, [if] I miss a line, I miss a path, I miss a way of calling a class method, etc, it corrects it on the spot and automatically shows the button: ‘Do you want me to correct it and run the cell again?’.” – P6 [Data Engineer].
Saving developer time with LLMs. Most of the participants (i.e., P1-P18, P20, P22) also describe how LLMs contribute to time-saving for simple, repetitive, or tedious tasks. The perceived saving in developers’ time happens due to the reduced effort provided by LLMs when the developer hands over control to LLMs, e.g.: “In the past, there were several sites, Stack Overflow […] From the moment you go looking for your question until you find it, you’ve already spent a lot of time.” – P6 [Data Engineer].
Gaining learning opportunities. Fourteen participants (i.e., P1-P4, P6-P7, P12, P14, P15, P17, P18, P20, P22) mentioned the potential of LLMs as an educational tool for software developers, as well as the learning opportunities from adopting LLMs. For instance, developers can search for the reason behind AI’s wrong suggestion, e.g.: “I try to see if the wrong answer was because I didn’t give enough information […] I’ll see if the prompt didn’t make sense” – P1 [Software Developer]. LLMs can provide alternative solutions that can reinforce developers’ solution arsenal, e.g.: “LLMs will give me an essentially different perspective. Maybe this is an alternative approach, a more concise way of doing things” – P17 [Data Analyst].
TABLE III: Summary of the Benefits of using LLMs at the Individual (Software Practitioner) Level.
SE Activity Software Development Tasks / How LLMs Impact? Why the Impact Happens?
Requirement Engineering & Software Design Task: Brainstorming
Benefit: [#1] Reducing effort, [#2] LLMs saving time, [#3] Gaining learning opportunities Reason:   LLMs Reducing $\text{Effort}^{\#2}$;  Mitigating $\text{interruptions}^{\#1,\#2}$; Handing over the Control to $\text{LLMs}^{\#1,\#2}$; $\text{Automation}^{\#1,\#2}$; Reliance on $\text{LLMs}^{\#1,\#2,\#3}$; Balanced control over $\text{LLMs}^{\#3}$; LLMs $\text{capabilities}^{\#3}$
Task: Requirement Documentation Generation
[…]
Tasks: Diagram Generation
[…]
Software Development & Software Quality Assurance Tasks: Concept Understanding
[…]
Tasks: Information Retrieval
[…]
Tasks: Code Understanding
[…]
Tasks: Code Generation; Code Translation; Code Documentation Generation; Code Comment Generation;  Unit Test Generation
[…]
Tasks: Test Case Identification, Test Data Generation
[…]
Tasks: Pull Request Generation
[…]
Software Maintenance Tasks: Debugging
Benefit: [#1] Reducing effort, [#2] LLMs saving time,[#3] Gaining learning opportunities,
[…]
Tasks: Code Review
[…]
Maintaining developer flow. Fifteen participants (i.e., P1, P4-P8, P12-P13, P15-P20, P22) presented different aspects that LLMs contribute to the software development flow. For instance, LLMs help developers to stay focused on the logic instead of typing when it autocompletes, e.g.: “So it helps me with […] reducing the time I’m actually writing and giving me more space to think about the flow in general.” – P4 [Software Developer].
Figure 3: Main Benefits of using LLMs at the Individual (Software Practitioner) Level.
Improving developer mental models. A few participants (i.e., P6, P10, P13-P14) mention how LLMs can aid developers in improving their code understanding, e.g.: “‘Why is it going on? Why is that method function [implemented in that way]?’” – P19 [Machine Learning Scientist].
Safe space. A few participants (i.e., P4, P7, P12, P15) mention how LLMs provide a safe space for sharing ideas and questions. For instance, software developers can ask as many questions as they need, e.g.: “I feel like senior people don’t have a lot of time on hand for trivial issues […] [but for LLMs] there is no judgment [while asking questions], because you can ask dumb questions to ChatGPT or any [other] LLMs. And you can ask it as many times as you want, and you can get examples […] You can get customised responses.” – P15 [Researcher]. This seems to be motivated by the desire to avoid disturbing colleagues’ work, while becoming self-reliant, e.g.: “Everyone’s very busy at work - the senior developers as well - you kind of - It’s not a set rule - but I feel I have a limited number of questions I can ask per day.” – P7 [BI Analyst]. In parallel, LLMs can offer certain levels of companionship experience, e.g.: “We can interact with LLMs like a person. And it’s good to have that kind of companionship as well, like you can talk to it and say: ‘Hey, what do you think about this?’” – P12 [Software Engineer], and empathy, e.g.: “There was one time when […] I imported some of the things Copilot sent me and tried them. When I went to Unity, it wouldn’t even run the project. […] [Copilot would reply]: ‘I’m sorry for your dissatisfaction, for your frustration, but let’s try to do it another way’.” – P4 [Software Developer].
3.1.2 Team Level
Mitigating interruptions. Fifteen participants (i.e., P1, P4-P6, P10, P12-P21) described how LLMs lead to fewer interruptions. Participants showed a preference to LLMs over disturbing colleagues, e.g.: “I’m always more comfortable using ChatGPT, Copilot. I do as little as possible to ask for help, so as not to disturb other people.” – P4 [Software Developer]. Participants also highlight the potential of LLMs supporting companies’ onboarding, which normally requires a lot of guidance from colleagues, e.g.: “If you’re kind of new to the team, you would feel more inclined to kind of just break it down through Copilot or ChatGPT first. And then afterwards, if you still don’t understand it, then kind of go ask an experienced coworker” – P14 [Data Engineer].
3.1.3 Organisation Level
Cost savings due to LLMs. A few participants (i.e., P7, P9, P15-P16, P20-P22) highlight how LLMs can contribute to reducing costs. For instance, using LLMs for debugging can decrease the time to find the problem, e.g.: “We are a small company, but we move at least one to 2 million dollars per week […] If there’s a production issue, then it’s like: ‘Oh, we need to solve this immediately’. And it [LLMs] saves a lot of time to find the bugs and to mitigate the problem. That [time spent searching for the problem] it costs, [and] you have put the money in somewhere else, where it should not be” – P12 [Software Engineer].
3.1.4 Societal Level
Foster Entrepreneurship. A few participants (i.e., P10 and P21) mention the potential of LLMs to encourage entrepreneurship. This occurs due to the great LLM prototyping capabilities, e.g.: “Prototyping becomes much faster for these [startup] companies […] Even a non-coder […] can create just the rundown prototype, not a full, scalable solution. But just to showcase sort of a working front-end” – P21 [Researcher Developer], and boost in individual productivity, reducing demand for big teams, e.g.: “It also makes it easier to create new companies […] because you can do more with fewer people. – P10 [Software Engineer]. At the same time, LLMs can provide basic information necessary to start a business, e.g.: “It can help you with the basics of marketing, accounting, and law, so it can help you tap into the basics of all these things, so you don’t need to recruit a team first.” – P21 [Researcher Developer].
RQ1. How do LLMs help software developers move forward? Our findings reveal benefits at the individual (nn=6), team (nn=1), organizational (nn=1), and societal (nn=1) levels. Specifically, LLMs assist software developers to sustaining cognitive flow, enhancing mental models, and offering a psychologically safe and continuous learning environment. LLMs can also reduce developer effort and save developer time.
3.2 RQ2. How do LLMs create an imbalance in software development?
3.2.1 Individual Level
Frictions in LLM-assisted software development. Most of the participants (i.e., P1-P3, P5-P19, P21, P22) mention how using LLMs may slow software developers. For instance, they may need to start a new conversation when LLMs get stuck in the same wrong suggestion, e.g.: “I usually started from scratch and gave him as much background and information as possible. Because when I continued the same conversation, it tended to continue the same mistake” – P6 [Data Engineer].
TABLE IV: Summary of the Challenges of using LLMs at the Individual (Software Practitioner) Level.
Rationale: The visualisation allows developers to create a mental model of code generated using LLMs.
  • Challenge: LLM-induced code quality risks. Rationale: The visualisation allows developers to review the workflow used in the code generated via LLMs.
  • Challenge: Risk of losing learning opportunities. Rationale: LLMs can suggest more appropriate architectures, overlooked by developers. This allows developers to improve their understanding of software design.
    Adopt self-hosted LLM solutions. Participants P9 and P22 run LLMs (i.e., Jan.AI, Llama, H2O Danube, and Mistral) locally via Ollama, Llama.cpp, and LM Studio as an approach to deal with privacy concerns in hosted LLM services.
    • Challenge: Issues with security & privacy
    Rationale: Employing self-hosted solutions eliminates concerns regarding the organisation’s data privacy.
  • Challenge: Risk of losing learning opportunities. Rationale: Deploying self-hosted solutions allows developers to have a better understanding of LLMs’ limitations.
    Integrate solutions, not just LLMs, into the workflow. Participant P9 defends that Copilot is an example of an AI-integrated solution, which comprehends GitHub Actions and GitHub – popular continuous integration and continuous delivery and code versioning platforms. Participant P22 incorporated an AI-assisted code review by connecting Ollama into their GitLab server.
    • Challenge: Developer flow interruptions through LLM interaction
    Rationale: Employing integrated LLM solutions allows for more contextualised and accurate assistance, reducing the need for human intervention.
    Combine different smaller models. After experimenting with H2O.ai [^3] https://h2o.ai, Participant P9 suggests using Small Language Models as a solid alternative to LLMs, e.g.: “In some parts of the solution, it’s possible to use smaller models very efficiently and diminish the cost of the overall solution” – P9 [AI Engineer].
    • Challenge: Cost of using LLMs
    Rationale: Employing smaller language models rather than large ones can help reduce computational, memory, and energy requirements [2].
    Prohibit LLMs from merging code into the main branch. Human in the loop is still required for accountability, e.g.: “If I’m putting up a pull request for my company, I need to understand that code. I need to have the ownership of it. Because if something fails, I won’t be blaming Cursor, or anyone else, this [error] will be on me to fix. Any code that’s generated by the LLMs, I only put this code in production [environment] if I actually understand that [code]” – P10 [Software Engineer].
    • Challenge: LLM-induced code quality risks
    Rationale: Implementing mandatory human review for main-branch commits can safeguard against errors in LLM-generated code [53].
  • Challenge: Reputational vulnerability through LLM delegation. Rationale: The credibility of developers and their organisations is preserved by safeguarding against errors from code generated via LLMs.
    Follow the established coding conventions. Participant P11 described how GitHub Copilot can get confused according to the used variables, e.g.: “When you have some kind of misleading or not super standard variable names, Copilot gets confused” – P11 [Web Developer].
    • Challenge: Frictions in human-LLM communication
    Rationale: By adhering to code conventions consistent with those found in the data used to train LLMs, developers can reduce the risk of miscommunication with these models.
  • Challenge: LLM-induced code quality risks. Rationale: By adhering to code convention best practices, code generated by LLMs is more likely to achieve satisfactory quality.
  • Challenge: Reputational vulnerability through LLM delegation. Rationale: By mitigating errors from code generated using LLMs, the credibility of developers and their organisations is preserved.
  • Challenge: Erosion of developer skills and agency. Rationale: By following code conventions, developers.
    Introduce LLM4SE usage policies. Four participants (18.1%) indicated in the pre-interview questionnaire that their organisations either do not have an AI policy or are not aware of one. Having a clear policy addressing the use of software development forms an essential foundation for effective IT governance.
    • Challenge: Issues with security & privacy
    Rationale: LLMs4SE policies that indicate the supported LLMs tools and how to share code with those tools may mitigate concerns regarding privacy.
  • Challenge: Issues with code license. Rationale: Defining how to handle code licensing in the LLM4SE policy can guide developers and help prevent potential legal issues in the future.
  • Challenge: Erosion of social trust. Rationale: Clarifying how LLMs manage developers’ prompts may help ease their concerns.
    Continuously upskilling in the latest SE practices and tools. Participants P1, P4, P10, P14, and P20 defend the importance of upskilling to not get behind in the market, e.g.: “I will always try to upskill myself in a way that I won’t get replaced by LLMs because the implementations are not perfect” – P20 [Software Engineer].
    • Challenge: Job market crisis
    Rationale: Keeping up-to-date with the latest SE best practices and tools, such as spec-driven development, may help developers to maintain relevance in this AI era.
    4 Discussion
    4.1 Implications for Practice
    We analyse the implications of adopting LLMs in SE by exploring both the benefits (RQ1) and challenges (RQ2) using Marshall McLuhan’s Tetrad of Media Effects [41]. This framework enables us to examine how new technologies can augment human capabilities through different aspects. The Tetrad consists of four fundamental questions: What does the technology enhance or intensify? What does it make obsolete or displaced? What does it retrieve or recover from the past? And what does it reverse into when taken to extremes?
    Enhance. This explores how adopting LLMs enhances or intensifies practices and processes in software development.
    • Cognitive debt: In section LABEL:subsubsec:incompleteModel, we highlight how adopting LLMs can deprive developers of a complete mental model of the code. According to Storey [52], this cognitive debt may silently build up over time and lead to developers losing control over their software. In our study, participants keep control over LLMs by (i) relying on them for code improvement rather than code generation and (ii) employing Test-Driven Development (TDD). Mathews et al. [39] found that employing TDD with models like GPT-4 and Llama 3 leads to a higher success rate in solving programming challenges.
    • Company onboarding: In section 3.1.2, we present the potential of LLMs being incorporated into the onboarding process. However, in the ongoing case study by Azanza et al. [3] at an IT consulting organisation, they found that using third-party LLMs can pose unacceptable data risks for companies. They explored open-source LLMs for onboarding. Additionally, organisations may seek agreement with LLM vendors for data protection.
    Obsolete. This aspect explores the obsolescence or displacement caused by adopting LLMs for software development.
    • Repetitive tasks: In section 3.1.1, we mention how LLMs reduce developers’ effort for simple tasks, such as boilerplate code generation and information retrieval. However, it is still essential to keep humans in the loop, at least as reviewers, since LLMs are based on stochastic models.
    Retrieve. This examines how far adopting LLMs makes it possible to retrieve or recover concepts, practices, or processes adopted in the past in SE.
    • Exploratory learning: In section 3.1.1, we highlight how participants use LLMs as a learning tool, augmenting the traditional problem-solving process. Using the chat mode, similar to Hallway talks, developers can ask questions, simulating hypothetical scenarios, getting an independent perspective. We suggest that other developers be added to the loop for sharing knowledge purposes.
    Reverse. This explains how adopting LLMs for software development, when pushed to its extreme, undergoes a reversal or transformation.
    • Skill atrophy: In section LABEL:subsubsec:skillAtrophy, we highlighted that adopting LLMs may hinder developers’ skills. While LLMs can provide time-saving for software development by reducing effort, that effort is also necessary to practice their skills. In our study, participants restrict themselves from overusing AI (e.g., turning off AI-based suggestions once in a while) and force themselves to use their coding muscles.
    • Environmental crisis: When it comes to environmental cost, surprisingly, only two participants (P16, P22) demonstrated concerns about energy consumption, overlooking water consumption [25]. Shi et al. [50] suggested that developers employ program-centric techniques, such as program pruning and grammar augmentation, to reduce the number of tokens.
    4.2 Implication for Research
    Based on the findings, we identified research gaps related to software practitioners adopting LLMs for Software Development.
    What factors influence developers’ decisions to delegate SE tasks to LLMs? Our interview participants described adopting different attitudes regarding the use of LLMs for SE tasks. We observed that task risk and developers’ motivation play an undeniable role in task delegation to LLMs. For instance, Participant P19 argues against relying on LLMs for test generation, warning that it could create a false sense of security. Regarding developers’ motivation, a lack of enthusiasm for certain tasks (e.g., updating documentation) among developers seems to encourage them to delegate such responsibilities. We suggest further studies to investigate other factors and how they interplay with task risk and developers’ motivation.
    How do LLMs empower developers during tight deadlines? Our data reveals a spectrum of behaviours: whereas some developers use LLMs to meet tight deadlines rapidly, others may inflate their effort estimations when motivation is low. Researchers should further examine the triggers for such inflation to better understand the long-term impact of AI assistance on software process metrics.
    How do LLMs affect social intelligence? Our findings show that using LLMs may mitigate interruptions at the team level. Although it may benefit in terms of productivity, the software development life cycle is entwined with “collaboration, human judgment, emotional interactions, and decision-making” [1, 28]. This is particularly important for manager positions, where social intelligence plays a crucial role in capitalising on interactions with customers and employers.
    How do LLMs affect developers’ intuition? According to Naur [44], “Software development in all its phases, and irrespective of the techniques employed in its pursuits, must and will always depend on intuition”. Intuition built from previous experiences acts as a safeguard mechanism, enabling software developers to navigate LLMs’ wrong suggestions. While Chen et al. [7] investigated the role of human intuition when using LLMs, the study regarding the effects of LLMs on human intuition remains unexplored.
    How does agentic software engineering contribute to developers’ burnout? Investigations into burnout in SE have been the focus of numerous studies [56]. In their systematic mapping study, Tulili et al. [56] identified the following causes for burnout: personality traits (e.g., neuroticism), work-related factors (e.g., job demand, job overload), communication practices, agile practices, and physiological factors. However, software developers’ burnout from adopting LLM tools is an emergent topic that requires attention from researchers. A recent large survey conducted by Feng et al. [17] found that job demands—such as organisational pressures and increased workload—are positively associated with burnout, while job resources related to generative AI adoption—such as autonomy and access to learning resources—are negatively associated with it. Beyond the adoption of generative AI, the rise of agentic software development can place a considerable burden on developers, who may need to supervise several agentic tools doing different software engineering tasks in parallel. Future research should investigate how this expanding supervisory burden interacts with established burnout drivers — and critically, what constitutes a sustainable threshold of concurrent agent supervision for software developers.
    How does agentic software engineering influence job title inflation within the tech industry? Title inflation occurs when employees are granted senior-level designations that exceed their actual duties or compensation [48]. In the era of agentic workflows, developers across all seniority levels are increasingly tasked with overseeing autonomous agents in addition to their traditional responsibilities. However, as these roles evolve to include agent orchestration, compensation structures often remain static. We propose that future studies investigate whether this gap between responsibility and reward is being bridged by inflated job titles and what impact this has on the long-term transparency of the software labour market.
    What are the most appropriate LLM maturity model stages for organisations? Our data shows varying degrees of LLM adoption, ranging from superficial tool access (e.g., GitHub Copilot for code completion) to deep integration within DevOps pipelines and code review processes. While the Quotient maturity framework [47] identifies five stages—ranging from ad hoc adoption to end-to-end autonomy—and suggests that stages 3 (standardised workflows) and 4 (supervised automation) are the most appropriate targets, these recommendations lack broad empirical support. We advocate for longitudinal studies to validate which maturity levels actually yield the highest return on investment and team stability in practice.
    Controversial use cases. Participants demonstrate different opinions regarding using LLMs for test generation, code generation, and debugging. LLMs are discouraged for test-related tasks due to providing the illusion of false assurance. Wang et al. [57] argue that using LLMs for integration tests may result in errors or unreliable results because those tests may exceed the capacity of the LLM to process and analyse. LLMs are also discouraged for code generation due to the harmful long-term effects towards software developers’ skills. Some participants also highlight that LLMs may provide general debugging guidance, since they cannot run the code.
    5 Related Work
    Figure 5: Comparison with related work focused on benefits and challenges. The colored bars represent the overlap with existing literature [5, 4, 35, 34], and empty bars indicate new findings.
    Barke et al. [5] conducted an observational and interview study with twenty software developers using GitHub Copilot on how they interact with GitHub Copilot. They employed traditional Grounded Theory analysis [20], identifying two interaction modes: acceleration mode and exploration mode. In acceleration mode, the developer has a clear understanding of the next required task; whereas developers explore GitHub Copilot autocomplete suggestions in exploration mode due to a lack of direction. According to Barke et al., the interactions in the acceleration mode are fast and do not disrupt the programmer’s flow. Our study findings corroborate their findings regarding LLMs maintaining development flow. Also aligned with our findings, they identified that LLMs can also disrupt the development flow when providing long suggestions.
    Banh et al. [4] conducted eighteen interviews with IT-related professionals (i.e., software developers, a product owner, and a scrum master) between August 2023 and January 2024. They aimed to understand how to integrate Generative AI in Software Engineering, exploring the opportunities and challenges related to adopting Generative AI tools. They applied the traditional Grounded Theory, resulting in the conceptual framework of generative AI integration in Software Engineering practices. Their findings, similar to ours, are: reduced development time, issues with intellectual property, underestimated overhead, developer empowerment, and code quality improvement via LLMs. However, our findings also include the contradictions involving the benefits and challenges, such as code degradation via LLMs, LLMs dragging tasks (see LABEL:subsubsec:disruptflow). Our results point out the necessity for a balanced approach to the adoption of LLMs, advocating for their use as a net positive in the software development process. Our recommendations in Section LABEL:subsec:recommendations contain practical suggestions contributing to achieve this.
    Liang et al. [35] employed traditional Grounded Theory [26] through twenty interviews and observation of fifteen software developers. We could not identify when interviews and observation were conducted. Their contribution includes an understanding of prompt programming practices. Similar to our study, they identified the potential of LLMs to assist developers in improving their mental model. Although their research focused on prompt programming aspects, they present their impact across different software development tasks similar to us (See Table 3.1.1 and LABEL:tab:disadvantageIndividualLevel). Our findings distinguish by showing not only the impact of software practitioners’ interaction with LLMs, but also how it reflects in others levels, such as team level (e.g., mitigating interruptions).
    Li et al., [34] conducted twenty-six interviews in three rounds with industry practitioners and 395 survey respondents. They employed socio-technical grounded theory [22, 23], which resulted in the theory of AI tool use and adoption in software engineering. Similar to our study, they also classified the impact at the individual and organisational level, such as fear of decreased skills and potential judgment of using LLMs, which are related to the sections hindering developers’ skills and damaging developers’ reputation from our study (See 3.2). On the other hand, our findings also comprehend the impact at the team and society level, such as losing mentorship opportunities due to LLMs, and LLMs fostering entrepreneurship.
    Although the literature presents studies exploring the impact of LLMs for software development, as illustrated in Figure 5, our investigation goes further by identifying best practices to balance the forward and backward impact of adopting LLMs.
    6 Evaluation
    6.1 Evaluating STGT4DA Application
    The evaluation of the STGT4DA application consists of credibility and rigour as key criteria [22, 23]. In terms of credibility, our section 2 provides details involving the recruitment process (social media and emailing), sampling method (purposive sampling followed by convenient sampling), how iterative and interleaved data occurred (three rounds of data collection and analysis), and how memos were written and applied (to guide the structure of emerging concepts and sub-categories, and as directions for future work). Concerning rigour, our section 2 also provides examples of basic coding and constant comparison (See Fig. 2) and embedded sanitised evidence (i.e., several interview quotes throughout Section 3).
    6.2 Evaluating STGT4DA Outcomes
    Findings from the application of STGT4DA should exhibit originality, relevance, and density [22, 23]. In terms of originality, we position our findings in comparison to related work in Section 5. The relevance of understanding the impact of LLMs on software practitioners is highlighted in Sections 1, 4.1, and 4.2. Regarding density, Section 3 condenses the sub-categories and concepts that compose the category “Impact on using LLMs”. We illustrate the depth of the findings through multiple levels of codes, concepts, and sub-categories, and the richness of findings through various aspects of the phenomenon, illustrated through numerous interview quotes.
    7 Threats to Validity
    We will discuss the study limitations using the Total Quality Framework (TQF) developed by [49], suitable for qualitative studies in software engineering [33, 21]. This framework is structured in the following aspects: credibility regarding data collection, analyzability of the data, transparency of reporting, and usefulness of the findings.
    Credibility. It refers to how comprehensively and accurately the data collection was performed [29]. Our study participants cover individuals from different demographics, such as gender, country of residence, years of professional experience, and role. At the same time, we acknowledge that future research could expand the participants’ pool by interviewing, for example, IT managers and project managers or collecting data from professional social media (e.g., LinkedIn and Reddit posts). Besides our challenges in recruiting participants - an arduous task described by [36] - we reached a suitable sampling size. We also acknowledge that changing the ATAI scale questions from 9 scale to 5 scale may impose threats to validity. However, we also defend that it may facilitate respondents to answer the questions.
    Analyzability. It refers to how comprehensively and accurately the data analysis was designed and executed. Although our qualitative analysis carries an intrinsic subjectivity, we mitigate this by conducting a pilot for the open coding, as well as conducting discussions over the emerging codes, concepts, sub-categories and categories between the authors.
    Transparency. It refers to how clear and complete this paper reports the aspects referring to credibility and analysability, supporting replication or transference to other contexts. We intend to present enough information. Additionally, we provide further information in the appendices and in the online supplementary material package [18].
    Usefulness. It refers to how useful the study’s contributions and findings are. Our findings can support software team leaders and IT managers in evaluating whether LLMs align with their context and needs, as well as guiding SE researchers for future research needs. We acknowledge that the fast-paced evolution of LLMs can quickly make some of our findings outdated. However, empirical studies help capture and document the evolving state of practice, especially as previous models become inaccessible. Further investigations can compare to our findings and report on the latest LLM capabilities.
    8 Conclusion
    In this paper, we present the software practitioners’ perspective on the impact of LLMs at the individual, team, organisation, and society levels. Our findings comprehend benefits and challenges of using LLMs for software development. At the individual level, most of the participants mention LLMs saving time and reducing effort as benefits, and frictions and overhead introduced by LLM-assisted software development as challenges. At the team level, most of the participants refer to LLMs as mitigating interruptions, but also reducing mentorship opportunities. At the organisation level, participants highlight the cost savings due to LLMs, but also security and privacy concerns. At the society level, participants mention LLMs promoting entrepreneurship, but also the erosion of social trust. The challenges identified create an imbalance in the use of LLMs. To address this issue, our findings also include actionable suggestions to help mitigate this imbalance and support developers as they navigate the integration of LLMs into software development.
    9 Acknowledgment
    Samuel is supported by a Faculty of IT PhD scholarship. We would like to express profound gratitude to all participants who took part in this study, Hashini Gunatilake, Humphrey Obie, and Marcos Medeiros, for the wonderful feedback on this research, and Nimmi Weeraddana and Haoyu Gao for helping during participant recruitment.